# Enhancing ground-state interaction strength of neutral atoms via Floquet stroboscopic dynamics

### v1

@@ V Implementation in a Rydberg Superatom
In principle, if the experimental techniques are sufficiently advanced to enable perfect Rydberg blockade and the impact of technical noise can be neglected, our protocol can be straightforwardly extended to systems with a large number of atoms [14, 109, 110, 111, 112]. In this idealized limit, the advantages of the scheme become even more pronounced, as the required number of driving cycles to prepare the target state decreases with an increasing atomic number. To gain a deeper understanding of the experimental realization of ground-state blockade in a Rydberg superatom, we reinterpret the mechanism in terms of interaction-enhanced single-excitation transfer. Under periodic driving, the system dynamics allow only a single atom to undergo repeated oscillations between $|e\rangle$ and $|r\rangle$, while multi-atom participation is inhibited. Due to the strong Rydberg blockade that forbids the simultaneous excitation of multiple atoms in $|r\rangle$, this process effectively suppresses the occurrence of multiple $|e\rangle$ excitations. In other words, the periodic protocol acts as if it reinforces the mutual exclusion among $|e\rangle$ excitations, ensuring that only a single $|e\rangle$ remains. For completeness, we also provide the approximate analytical form of the system evolution under simplifying assumptions in Appendix B.
As shown in Fig. 6, we numerically simulate the dynamics of the fidelity for the $|W_{\mathbb{N}}\rangle$ under different realizations of atom numbers $\mathbb{N}$ sampled from a predefined distribution. The fidelity is evaluated as $F(\mathbb{N}_{0})=\langle\psi_{\mathrm{target}}(\mathbb{N}_{0})|\rho_{\mathbb{N}_{0}}(t)|\psi_{\mathrm{target}}(\mathbb{N}_{0})\rangle$, where $|\psi_{\mathrm{target}}(\mathbb{N}_{0})\rangle=|W_{\mathbb{N}_{0}}\rangle$ denotes the $W$ state for the $\mathbb{N}_{0}$ atoms. This allows us to assess the robustness of the protocol against atom-number fluctuations. The pulse duration $t_{a}=\pi/(\sqrt{\mathbb{N}}N\omega)$ is designed for the reference atom number $\mathbb{N}=500$. The same control sequence, with this fixed pulse duration, is applied to all realizations with different actual atom numbers $\mathbb{N}_{0}$. Since the pulse duration is fixed, this deviation effectively manifests itself as a fluctuation in the effective evolution time rather than a change in the control protocol itself. The results demonstrate that even in the presence of $20\%$ fluctuations in the atom number around $\mathbb{N}=500$, the system still converges to the single-excitation state of the ground state with a high fidelity of $98\%$ after several driving cycles, highlighting the strong robustness of the scheme against atom number fluctuations.
@@ VI Corresponding Energy-Level Diagram and Transitions
In the experiment, we can find the atomic configuration required for our scheme in the 87Rb atom, as shown in Fig. 7. The ground states $|g\rangle$ and $|e\rangle$ correspond to the atomic levels $|5S_{1/2},F=1,m_{F}=0\rangle$ and $|5S_{1/2},F=2,m_{F}=0\rangle$, the Rydberg state is $|nS_{1/2},m_{j}=1/2\rangle$, $|5P_{3/2},F=2,m_{F}=1\rangle$ is an intermediate state that can be adiabatic eliminated under conditions $\Delta_{a}\gg\{\Omega_{1},\Omega_{2}\}$ and $\Delta_{p}\gg\{\Omega_{3},\Omega_{4}\}$, leading to a direct coupling between $|g\rangle$ and $|e\rangle$ with two-photon Rabi frequency $\omega$ and $|e\rangle$ and $|r\rangle$ with two-photon Rabi frequency $\Omega$. Ignoring the Stark shift, we can obtain the effective Hamiltonian of these two-photon processes as
@@
$H_{b}^{\mathrm{eff}}=\frac{\Omega_{3}\Omega_{4}}{4\Delta_{p}}|r\rangle\langle e|+\mathrm{H.c.}$ (8)
The corresponding Rabi frequencies used in the text read $\omega=\Omega_{1}\Omega_{2}/2\Delta_{a}$ and $\Omega=\Omega_{3}\Omega_{4}/2\Delta_{p}$. The Raman transition between the ground states $|g\rangle$ and $|e\rangle$ is driven by the use of two 780-nm laser beams with $\sigma^{+}$-polarized. The Rydberg excitation uses $\sigma^{+}$-polarized 780-nm and $\sigma^{-}$-polarized 480-nm beams tuned for excitation of the Rydberg state of $nS_{1/2}$. A wide range of experimentally accessible parameters can be used in our scheme. In particular, the Raman coupling between the ground states can be tuned over a broad interval, and values of the effective Rabi frequency $\omega/2\pi$ in the range of 0.530.5–3 MHz are readily achievable with standard optical configurations [52, 113, 114, 115, 116, 117]. Higher Rabi frequencies for the ground-to-Rydberg excitation are generally favorable, and current experimental techniques can provide sufficiently strong driving for our scheme [103, 71, 118, 51, 119]. For reference, Ref. [52] demonstrated a two-photon Raman coupling between the states $|5S_{1/2},F=1,m_{F}=0\rangle$ and $|5S_{1/2},F=2,m_{F}=0\rangle$ with an effective Rabi frequency of $\omega/2\pi=1.36$ MHz using an intermediate-state detuning of $\Delta_{a}/2\pi=-41$ GHz. In addition, Ref. [71] achieved a two-photon Rydberg excitation with intermediate-state detuning $\Delta_{p}/2\pi=7.8$ GHz and a Rabi frequency of $\Omega/2\pi=4.6$ MHz. The decay rate of the Rydberg state $|nS_{1/2},m_{j}=1/2\rangle$ is $\gamma\approx 2\pi\times 1$ kHz for the room-temperature (300 K) case. In the single-photon generation scheme, we designate the state $|5P_{3/2},F=2,m_{F}=1\rangle$ as the controlled excited state, which is coupled to the excited state $|5S_{1/2},F=2,m_{F}=0\rangle$ via a $\sigma^{+}$-polarized 780-nm. Furthermore, the scheme does not depend on the strength of van der Waals interactions. From this perspective, the proposed mechanism is adaptable to a wide variety of parameters.
@@ VII Conclusion
We have proposed a Floquet-engineered scheme to enhance the effective interaction between ground-state neutral atoms in a Rydberg atomic ensemble. By periodically driving the system with a sequence of ground-state couplings followed by a $4\pi$ pulse on the ground–Rydberg transition, the dynamics are confined to the single-excitation manifold through stroboscopic evolution. As a result, the system evolves from the collective ground state $|G\rangle$ to the state $|W_{\mathbb{N}}\rangle$, while double excitations remain strongly suppressed. This behavior reflects an effective ground-state blockade induced by Floquet modulation. Unlike conventional schemes that rely on strong Rydberg-Rydberg interactions and are therefore limited to the blockade radius, our approach remains effective even outside the conventional Rydberg blockade radius. In addition, to assess its feasibility under realistic conditions, we examine the impact of relevant experimental imperfections, including the spontaneous emission of the Rydberg state, laser phase noise, Rabi amplitude noise, and laser detuning errors. The results show that the scheme remains highly robust against these errors and can tolerate thermal fluctuations in atomic positions, making it readily implementable under realistic experimental conditions. Building on this mechanism, we further examine its application to single-photon sources. The simulated equal-time second-order correlation function indicates that the scheme can theoretically achieve a high-efficiency single-photon source. Finally, we explore the implementation of our scheme in a Rydberg superatom, demonstrating that shortening the Floquet driving period enables practical realization within a single Rydberg superatom. In summary, our scheme realizes robust quantum state preparation by effectively enhancing ground-state interactions in neutral atoms. This mechanism can be directly applied to deterministic single-photon generation, providing a versatile and high-fidelity platform for quantum technologies.
@@ Appendix B The Zeno Hamiltonian of two atoms and dynamics of multi-atomic symmetric state
$|P_{\mathbb{N}}\rangle$ $=\frac{1}{\sqrt{C_{\mathbb{N}}^{2}}}\sum_{i<j}|g_{1}\cdots e_{i}\cdots e_{j}\cdots g_{N}\rangle\quad$
@@
The target state $\big|\psi_{\mathrm{target}}(\mathbb{N}_{0})\big\rangle$ is chosen as the single-excitation state

### v2

@@ V Implementation in a Rydberg Superatom
In principle, if the experimental techniques are sufficiently advanced to enable perfect Rydberg blockade and the impact of technical noise can be neglected, our protocol can be straightforwardly extended to systems with a large number of atoms [15, 115, 116, 117, 118]. In this idealized limit, the advantages of the scheme become even more pronounced, as the required number of driving cycles to prepare the target state decreases with increasing atomic number. To gain a deeper understanding of the experimental realization of ground-state blockade in a Rydberg superatom, we reinterpret the mechanism in terms of interaction-enhanced single-excitation transfer. Under periodic driving, the system dynamics allow only a single atom to undergo repeated oscillations between $|e\rangle$ and $|r\rangle$, while multi-atom participation is inhibited. Due to the strong Rydberg blockade that forbids the simultaneous excitation of multiple atoms in $|r\rangle$, this process effectively suppresses the occurrence of multiple $|e\rangle$ excitations. In other words, the periodic protocol acts as if it reinforces the mutual exclusion among $|e\rangle$ excitations, ensuring that only a single $|e\rangle$ remains. For completeness, we also provide the approximate analytical form of the system evolution under simplifying assumptions in Appendix B.
As shown in Fig. 6, we numerically simulate the dynamics of the fidelity for the $|W_{\mathbb{N}}\rangle$ under different realizations of atom numbers $\mathbb{N}$ sampled from a predefined distribution. The fidelity is evaluated as $F(\mathbb{N}_{0})=\langle\psi_{\mathrm{target}}(\mathbb{N}_{0})|\rho_{\mathbb{N}_{0}}(t)|\psi_{\mathrm{target}}(\mathbb{N}_{0})\rangle$, where $|\psi_{\mathrm{target}}(\mathbb{N}_{0})\rangle=|W_{\mathbb{N}_{0}}\rangle$ corresponds to the $W$ state for the $\mathbb{N}_{0}$ atoms. Here, $\mathbb{N}_{0}$ represents the actual atom number in a given experimental realization, which may deviate from the nominal number $\mathbb{N}$ due to atom loss. This allows us to assess the robustness of the protocol against atom-number fluctuations. The pulse duration $t_{a}=\pi/(\sqrt{\mathbb{N}}N\omega)$ is designed for the reference atom number $\mathbb{N}=500$. The same control sequence, with this fixed pulse duration, is applied to all realizations with different actual atom numbers $\mathbb{N}_{0}$. Since the pulse duration is fixed, this deviation effectively manifests itself as a fluctuation in the effective evolution time rather than a change in the control protocol itself. The results demonstrate that even in the presence of $20\%$ fluctuations in atom number around $\mathbb{N}=500$, the system still converges to the symmetric collective state of the ground state with a high fidelity of $98\%$ after several driving cycles, highlighting the strong robustness of the scheme against atom number fluctuations.
@@ VI Corresponding Energy-Level Diagram and Transitions
In the experiment, we can find the atomic configuration required for our scheme in the 87Rb atom, as shown in Fig. 7. The ground states $|g\rangle$ and $|e\rangle$ correspond to the atomic levels $|5S_{1/2},F=1,m_{F}=0\rangle$ and $|5S_{1/2},F=2,m_{F}=0\rangle$, the Rydberg state is $|nS_{1/2},m_{j}=1/2\rangle$, $|5P_{3/2},F=2,m_{F}=1\rangle$ is an intermediate state that can be adiabatic eliminated under conditions $\Delta_{a}\gg\{\Omega_{1},\Omega_{2}\}$ and $\Delta_{p}\gg\{\Omega_{780},\Omega_{480}\}$, leading to a direct coupling between $|g\rangle$ and $|e\rangle$ with two-photon Rabi frequency $\omega$ and $|e\rangle$ and $|r\rangle$ with two-photon Rabi frequency $\Omega$. Ignoring the Stark shift, we can obtain the effective Hamiltonian of these two-photon processes as
@@
$H_{b}^{\mathrm{eff}}$ $=\frac{\Omega_{780}\Omega_{480}}{4\Delta_{p}}|r\rangle\langle e|+\mathrm{H.c.}$ (9)
The corresponding Rabi frequencies used in the text read $\omega=\Omega_{1}\Omega_{2}/2\Delta_{a}$ and $\Omega=\Omega_{780}\Omega_{480}/2\Delta_{p}$. The Raman transition between the ground states $|g\rangle$ and $|e\rangle$ is driven by the use of two 780-nm laser beams with $\sigma^{+}$-polarized. The Rydberg excitation uses $\sigma^{+}$-polarized 780-nm and $\sigma^{-}$-polarized 480-nm beams tuned for excitation of the Rydberg state of $nS_{1/2}$. A wide range of experimentally accessible parameters can be used in our scheme. In particular, the Raman coupling between the ground states can be tuned over a broad interval, and values of the effective Rabi frequency $\omega/2\pi$ in the range of 0.530.5–3 MHz are readily achievable with standard optical configurations [53, 119, 120, 121, 122, 123]. Higher Rabi frequencies for the ground-to-Rydberg excitation are generally favorable, and current experimental techniques can provide sufficiently strong driving for our scheme [108, 73, 124, 52, 125]. For reference, Ref. [53] demonstrated a two-photon Raman coupling between the states $|5S_{1/2},F=1,m_{F}=0\rangle$ and $|5S_{1/2},F=2,m_{F}=0\rangle$ with an effective Rabi frequency of $\omega/2\pi=1.36$ MHz using an intermediate-state detuning of $\Delta_{a}/2\pi=-41$ GHz. In addition, Ref. [73] achieved a two-photon Rydberg excitation with intermediate-state detuning $\Delta_{p}/2\pi=7.8$ GHz and a Rabi frequency of $\Omega/2\pi=4.6$ MHz. In the single-photon generation scheme, we designate the state $|5P_{3/2},F=2,m_{F}=1\rangle$ as the controlled excited state, which is coupled to the excited state $|5S_{1/2},F=2,m_{F}=0\rangle$ via a $\sigma^{+}$-polarized 780-nm. Furthermore, the scheme does not depend on the strength of vdW interactions. From this perspective, the proposed mechanism is adaptable to a wide variety of parameters.
@@ VII Conclusion
We have proposed a Floquet-engineered scheme to enhance the effective interaction between ground-state neutral atoms in a Rydberg atomic ensemble. By periodically driving the system with a sequence of ground-state couplings followed by a $4\pi$ pulse on the ground–Rydberg transition, the dynamics are confined to the single-excitation manifold through stroboscopic evolution. As a result, the system evolves from the collective ground state $|G\rangle$ to the state $|W_{\mathbb{N}}\rangle$, while double excitations remain strongly suppressed. This behavior reflects an effective ground-state blockade induced by Floquet modulation. Unlike conventional schemes that rely on strong Rydberg-Rydberg interactions and are therefore limited to the blockade radius, our approach remains effective even outside the conventional Rydberg blockade radius. In addition, we evaluate the feasibility of the proposed scheme under realistic experimental conditions by systematically incorporating major sources of imperfections, including Rydberg-state spontaneous emission and blackbody-radiation-induced decay, laser phase noise, Rabi amplitude fluctuations, detuning errors, as well as Doppler-induced dephasing arising from finite atomic temperature. Our results demonstrate that the scheme exhibits strong robustness against these noise sources and remains resilient to thermal motion and ensemble inhomogeneity, indicating that it is well suited for implementation in current neutral-atom platforms. We note that additional noise sources specific to particular experimental implementations may also contribute to decoherence but are beyond the scope of the present theoretical model. Building on this mechanism, we further examine its application to single-photon sources. The simulated equal-time second-order correlation function indicates that the scheme can theoretically achieve a high-efficiency single-photon source. Finally, we explore the implementation of our scheme in a Rydberg superatom, demonstrating that shortening the Floquet driving period enables practical realization within a single Rydberg superatom. In summary, our scheme realizes robust quantum state preparation by effectively enhancing ground-state interactions in neutral atoms. This mechanism can be directly applied to deterministic single-photon generation, providing a versatile and high-fidelity platform for quantum technologies.
@@ Appendix B The Zeno Hamiltonian of two atoms and dynamics of multi-atomic symmetric state
$|P_{\mathbb{N}}\rangle$ $=\frac{1}{\sqrt{C_{\mathbb{N}}^{2}}}\sum_{i<j}|g_{1}\cdots e_{i}\cdots e_{j}\cdots g_{\mathbb{N}}\rangle\quad$
@@
The target state $\big|\psi_{\mathrm{target}}(\mathbb{N}_{0})\big\rangle$ is chosen as the symmetric collective state